### v1

# Unbounded entropy production and violent fragmentation in long-range interacting super-Tonks-Girardeau systems

Abstract
We study the non-equilibrium dynamics of a one-dimensional Bose gas with long-range interactions that decay as $(\frac{1}{r^{\alpha}})$ $(0.5<\alpha<4.0$). We investigate the system when the interactions are suddenly switched from the Tonks-Girardeau (TG) to the super-Tonks-Girardeau (sTG) limit. We find that relaxation is achieved through a complex intermediate dynamics demonstrated by violent fragmentation and chaotic delocalization. We establish that the sTG gas exhibits classical gaseous characteristics and an asymptotic state associated with unbounded entropy production. The phase diagram shows an exponential boundary between the nonthermal (quantum) gas and the thermal (classical) gas. We show the universality of the dynamics by also presenting results for spinless fermions. Weaker quench protocols give a certain degree of control over the relaxation process and induce some intermediate quasi-prethermal states. Our study showcases the complex relaxation behavior of tunable long-range interacting systems that could be engineered in state-of-the-art experiments, e.g. in trapped ions or Rydberg atoms.
entropy, long-range interactions, crystal states, super-Tonks-Girardeau quenches
I Introduction
Quantum systems with long-range interactions that decay as $\frac{1}{r^{\alpha}}$ have recently received increasing interest due to rapid developments of experimental techniques for controlling and manipulating atomic, molecular and optical (AMO) systems [1]. Long-range quantum systems are currently being realized in several experimental platforms such as Rydberg atoms [2], trapped ions [3, 4], polar molecules [5], and dipolar quantum gases [6]. Furthermore, long-range interactions with a tunable exponent $\alpha$ can be realized in Paul traps [7, 8]. Out of all these platforms, trapped ions present the most unique possibility of exploring long-range interactions with a decay constant $\alpha$ tunable between 0 and 3. Moreover, recent experimental setups with ion traps enable the realization of spin chains with controlled long-range interactions [9, 10, 11].
The dynamical properties of long-range interacting systems significantly differ from those of short-range interacting ones in many aspects [12, 13, 14, 15]. Depending on the exponent $\alpha$, long-range interactions can exhibit novel features. This is especially true in the very long-ranged regime when $\alpha<d$ ($d$ being the dimension of the system). These include long-lived metastable states [16, 17], supersolids and dipolar supersolids [18, 19, 20, 21, 22, 23, 24, 25, 26, 27, 28], dynamical phenomena such as time crystals [29, 30, 31, 32] and various Floquet phases [33, 34, 35, 36, 37].
The nonequilibrium dynamics and (pre)thermalization of isolated quantum systems with short-range interactions is rather well understood and has been extensively studied [38, 39, 40, 41, 42, 43, 40]. On the other hand, long-range interactions introduce more intriguing complexities which are poorly understood [44, 45, 46, 1]. For example, short-range interacting systems typically give rise to a single time scale of thermalization, while for strong long-range interactions thermalization can occur in multiple time steps [46].
The simplest protocol for probing such nonequilibrium dynamics is a quantum quench, which consists of preparing the initial setup in the ground state of a given Hamiltonian and then suddenly quench one of its parameters to a different value at time zero. Thermalization [47] happens when a local observable approaches a statistical value or equilibrium [48, 49]. However, depending on the quench parameter, an initial quantum system can exhibit a two step relaxation process instead. The local observable first settles to a quasi-stationary value, and only at a later stage reaches thermalization. This phenomenon is termed prethermalization [50]. Prethermalization happens when there is a clear separation in relevant time scales with different physical origins, which typically depend on individual systems [51].
Prethermalization was directly observed in a one-dimensional (1D) Bose gas of 87Rb atoms [52]. In this experiment, the initial-phase fluctuating 1D Bose gas is split into two uncoupled gases in a double well potential with almost identical phase difference. The phase coherence between the two gases is further studied as a function of the evolution time. After some initial rapid evolution, a thermal-like steady state is approached – which is associated with prethermalization – before the system eventually equilibrates.
Exact theoretical studies of (pre)thermalization in long-range interacting systems for both bosons and fermions are very limited [53, 13]. Study on the quench dynamics with long-range interacting spin chains so far concluded that prehermalization occurs for $10<\alpha<$ 1 [46]. In another theoretical study, utilizing spatial inhomogeneity, three different relaxation processes were observed: prethermalization only, prethermalization followed by thermalization, and thermalization only [54]. Prethermalization has also captured the attention in spinless fermions and in the context of fermionic Hubbard models [55, 56, 57]. Although significant amount of work both theoretical and experimental has been devoted to understanding the prethermalization process, its physical origin is still elusive. It is also not clear whether it can be related to the quasi-integrability of the system. For bosons, theoretical studies so far addressed the cases of quasi-BEC regimes [51] and the case of strongly interacting dipolar bosons remains rather unexplored.
Our main motivation is to establish an all-in-one out-of-equilibrium dynamics for the generalized long-range interactions that can be observed in AMO experiments. To that end, the present work focuses on generalized long-range interacting bosons and fermions in the strongly interacting regime given by the Tonks-Girardeau (TG) limit. The TG gas is one of the most remarkable examples of quantum correlation amplification arising from a strong interplay between interatomic interactions and one-dimensional (1D) confinement. In the TG limit, bosons with strong repulsive interactions escape their spatial overlap to minimize their interaction energy and thereby acquire fermionic properties [58], which have been experimentally observed [59]. From a TG gas, it is possible to stabilize another highly excited and correlated state by quenching the interactions to the strongly attractive regime instead. This state is called a super-Tonks-Girardeau gas (sTG) [60, 61]. Intuitively, strong attractions should lead the system to quickly cluster. Remarkably, instead, the TG arrangement is preserved and the many-body phase relaxes to a metastable state, as verified experimentally in the work of Haller et al. [62]. In the experiment, a magnetically induced Feshbach resonance [63] is employed to control the 1D scattering properties by tuning the coupling constant of the 1D $\delta$-function contact potential $(g_{1D})$ [64]. The TG to sTG quench is achieved when $g_{1D}$ is switched from +$\infty$+ ∞ to $\infty$- ∞, which leads to the excited gas phase.
Theoretically, the sTG quench problem has been addressed by solving the dynamics of a Lieb-Liniger Bose gas [65], but paradigmatic questions regarding relaxation dynamics are still not clarified: i) What is the underlying microscopic dynamics that transforms the TG phase into the sTG phase? ii) How does the non equilibrium dynamics change when longer-ranged interactions are employed? iii) How can the excited gas phase be characterized from a many-body perspective? iv) Does equilibration happen in a single or dual time scale? v) Is it possible to control the relaxation process in the quench dynamics? Motivated by the recent experiments with magnetic atoms [66, 67], polar molecules [68, 69] and the flexibility to tune the long-range interactions [70], we address all these questions by studying sTG quenches for long-ranged systems with varying exponent $\alpha$.
We solve the time-dependent Schrödinger equation (TDSE) for the sTG quench process numerically by employing the MultiConfiguration Time-Dependent Hartree method for bosons (MCTDHB) [71, 72, 73, 74] implemented in the software MCTDH-X [74, 75, 76, 77, 78, 79, 80]. MCTDH-X has been a very prolific tool for studying the dynamics of long-range interacting quantum gases in the past [81, 82, 83, 84, 85, 86, 87, 88, 89, 90, 91, 92, 93, 94, 95, 96, 97, 98], as it enables us to evaluate the full time evolution of time-dependent processes in a very efficient way.
We follow the experimental protocol of Haller [62] and consider generic power-law interactions with $40.5\leq\alpha<$ 4.
For each value of $\alpha$, we first prepare $N=5$ strongly repulsive bosons in the ground state of a harmonic trap, obtaining a crystal state. Then, at time $t$=0, we suddenly quench their interactions to the strongly attractive regime and monitor the particle dynamics. We find that the time evolution exhibits violent dynamical fragmentation, unbounded entropy production, and complete delocalization in configuration space. These three features suggest that the highly excited sTG phase behaves as a classical gas when starting from a crystal state exhibiting long-range interactions. We also perform analogous simulations with fermions, observing a very similar behavior which indicates the universality of our results in the dynamics of non-local interacting systems. For both bosons and fermions, the system eventually relaxes to a fully occupied configuration space for the entire range of $\alpha$. We do not find any signatures of any intermediate time scale in the relaxation process. However, we observe an exponential boundary in the entropy evolution between relaxed and non-relaxed states, indicating a progressive slowdown in the relaxation process at increasing longer-ranged interactions (lower value of $\alpha$). To understand the process of relaxation, the system is also studied for quenches to other weakly attractive regimes. We observe that the relaxation process is less violent, slows down and begins to exhibit very short lived quasi-equilibrium states when $\alpha<1$.
The present study sheds light on dynamical universal behavior of highly non-local interacting systems, which directly relates to the large amount of research in state-of-the-art experiments in the AMO systems. We envision a rich interplay between long-range interactions and post quench dynamics both for ultracold bosons and fermions. We believe that our in-depth analysis can be exploited to understand the richness of these versatile quantum long-range systems.
The paper is structured as follows. In Sec. II, we discuss the Hamiltonian and theoretical approach. In Sec. III, we introduce quantities of interest. Sec. IV discusses the quantum to classical behavior through the analysis of three key quantities: orbital occupation, Fock space occupation, and entropy measures. Sec. V illustrates the underlying relaxation process both for bosons and fermions. Sec. VI presents how to control the quench mechanism to achieve almost prethermal states. We conclude our paper in Sec. VII. Appendix A presents an overview of the initial states and Appendix B considers a stronger quench dynamics.
II Model and theoretical approach
The time evolution of $N$ interacting bosons is governed by the TDSE (here and henceforth, we set $\hbar=1$)
$i\frac{\partial\psi}{\partial t}=\hat{H}\psi,$ (1)
where the total Hamiltonian for the system is
$\hat{H}(x_{1},\dots,x_{N})=\sum_{i=1}^{N}\hat{h}(x_{i})+\sum_{i<j=1}^{N}\hat{W}(x_{i}-x_{j})$ (2)
The Hamiltonian $\hat{H}$ is expressed in dimensionless units  [^1] The Hamiltonian $\hat{H}$ is obtained by dividing the dimensionful Hamiltonian by $\frac{\hbar^{2}}{mL^{2}}$, with $m$ the mass of the particles and $L$ an arbitrary length scale, which we choose to be the point at which the harmonic trap equals $\frac{1}{2}$.. The operator $\hat{h}(x)=\hat{T}(x)+\hat{V}(x)$ is the one-body Hamiltonian containing the kinetic energy $\hat{T}(x)=-\frac{1}{2}\hat{\partial}_{x}^{2}$ and the external potential $\hat{V}(x)=\frac{1}{2}x^{2}$. The operator $\hat{W}(x_{i}-x_{j})$ describes the two-body interaction between particles at positions $x_{i}$ and $x_{j}$ (which can be either attractive or repulsive).
The MCTDH ansatz for the bosonic many-body wave function is a linear combination of time dependent permanents constructed over $M$ single-particle wave functions called orbitals  [^2] The ansatz for fermionic many-body wave functions is equivalent, but Slater determinants replace the permanents. We redirect to Ref. [75] for further details.:
$|\psi(t)\rangle=\sum_{\bar{n}}C_{\vec{n}}(t)|\vec{n};t\rangle.$ (3)
The vector $\vec{n}=(n_{1},n_{2},\dots,n_{M})$ represents the occupation of each orbital with the constraint that $n_{1}+n_{2}+\dots+n_{M}=N$, which ensures the preservation of the total number of particles. Distributing $N$ bosons over $M$ time dependent orbitals, the number of configurations becomes
$N_{\mathrm{conf}}=\left(\begin{array}[]{c}N+M-1\\ N\end{array}\right).$ (4)
It is important to emphasize that in the ansatz (3), both the expansion coefficients $C_{\vec{n}}(t)|\vec{n};t\rangle$ and the orbitals that build up the permanents $|\vec{n};t\rangle$ are time-dependent, fully variationally optimized quantities. This time-adaptive basis allows the sampled Hilbert space to dynamically follow the motion under any quench dynamics.
In the limit $M\rightarrow\infty$, the representation of Eq. 3 is exact. However, in practice, the size of the Hilbert space has to be truncated by using a finite value of $M$ during computation. The time-dependent permanents assure that a given degree of accuracy is reached with a much shorter expansion compared to the time independent basis used in exact diagonalization.
The occupation of the different orbitals offers a measure for the many-body state dynamical fragmentation in the nonequilibrium dynamics of the quantum quench. More precisely, we define fragmentation from the natural occupations $n_{i}$, i.e. the population of the natural orbitals, which are the eigenvalues of the reduced one-body density matrix $\rho^{(1)}(x,x^{\prime};t)=\langle\psi(t)|\hat{\psi}^{\dagger}(x^{\prime})\hat{\psi}(x)|\psi(t)\rangle$, i.e.
$\rho^{(1)}(x,x^{\prime},t)=\sum_{i}n_{i}\phi_{i}^{*}(x^{\prime})\phi_{i}^{*}(x)$ (5)
as a spectral decomposition. In the expression for the reduced one-body density matrix, $\hat{\psi}^{(\dagger)}(x)$ are field operators for the creation and annihilation of one particle at position $x$. For a non-fragmented or condensed state, the occupation of the lowest natural orbital is close to unity. However, for a fragmented state, several natural orbitals may contribute significantly. In particular, for bosons, (lack of) fragmentation illustrates how close the many-body state is to a global superfluid that can be described by a single wave function.
III Quantities of interest
To study statistical relaxation, the Shannon information entropy (SIE) $S^{\mathrm{info}}(t)$ is the ideal quantity [101]. The SIE of the one-body density in coordinate space is defined as $S^{\mathrm{info}}(t)=-\int{\mathrm{d}x\rho^{(1)}(x,t)\ln\left[{\rho^{(1)}(x,t)}\right]}$. The SIE is a measure of localization/delocalization of the corresponding density. However, it is insensitive to the correlation present in $|\psi(t)\rangle$. Considering the MCTDHB ansatz (3), then, we define a few alternative entropy measures [83]. The coefficient Shannon information entropy is defined as
$S_{C}(t)=-\sum_{\vec{n}}|C_{\vec{n}}(t)|^{2}\ln|C_{\vec{n}}(t)|^{2}.$ (6)
It characterizes the distribution of the state $|\psi(t)\rangle$ in the Fock space. $S_{C}=0$ indicates a single-configuration state or a mean-field state. When $|\psi(t)\rangle$ spreads over the available space determined by the configurations, several expansion coefficients become nonzero and $S_{C}$ gradually increases. When the $N$-body Fock space is completely populated, $S_{C}$ saturates to the maximal value. Thus, in a quench procedure, $S_{C}(t)$ measures the degree of Fock space delocalization with time.
The occupation Shannon information entropy is defined as
$S_{n}(t)=-\sum_{i}\bar{n}_{i}(t)[\ln\bar{n}_{i}(t)],$ (7)
where $\bar{n}_{i}(t)=\frac{n_{i}(t)}{N}$ are natural occupations normalized with the particle number. In the single-orbital mean-field theory, as the reduced density matrix has only a single eigenvalue, $S_{n}=0$ and the corresponding many-body state is non-fragmented. When multiple significant orbitals contribute, the corresponding many-body state becomes fragmented. Thus, in a relaxation process, $S_{n}(t)$ initially starts with a minimum value (which may not be zero), increases and finally saturates to a maximal value. Therefore, $S_{n}(t)$ conveys the emergence and degree of fragmentation.
Finally, we can also study entropy measures obtained from the coefficients of the many-body decomposition. The $N$-body coefficient entropy $S_{C}^{N}(t)$ is defined as a hybrid version of the coefficient Shannon information entropy,
$S_{C}^{N}(t)=-\sum_{\vec{n},\vec{n^{\prime}}}|C_{\vec{n}}(t)|^{2}\ln|C_{\vec{n^{\prime}}}(t)|^{2}.$ (8)
This quantity conveys information about delocalization in Fock space, but considering all possible pairings of configurations. A related quantity is the coefficient inverse participation ratio. This is another measure to detect the effective number of basis states participating in the time evolution of the many-body state and is generally utilized to understand irregular dynamics. It is defined as
$I_{C}(t)=\frac{1}{\sum_{\vec{n}}|C_{\vec{n}}(t)|^{4}}.$ (9)
When $I_{C}(t)$ is small, the corresponding many-body state is close to the mean-field state and the corresponding dynamics is regular. Whereas large $I_{C}(t)$ indicates that the corresponding dynamics is irregular or chaotic.
IV Quantum to classical behavior
We now present the results of our quench dynamics calculations. We consider generalized long-range (power law) interactions determined by the exponent $\alpha$:
$\hat{W}(x_{i}-x_{j})=\frac{g_{d}}{|x_{i}-x_{j}|^{\alpha}+\epsilon_{\alpha}}.$ (10)
The parameter $g_{d}$ controls the strength of the long-range interactions. In the current platform of generalized AMO systems, $\alpha$ is scaled according to the type of interaction and generally the ratio $\frac{\alpha}{d}$ measures how strong the interaction is. For our present study in 1D system, $\alpha$ simply measures the range of interaction considering both weak long-range and strong long-range interactions. In general, $\alpha=1$ appears for gravitational or Coulomb interactions, $\alpha=3$ describes a dipolar quantum gas, and $\alpha=0$ corresponds to interaction mediated by cavity photons (infinite ranged). For $\alpha=3$ (dipole-dipole interactions), $g_{d}$ could be expressed as $g_{d}=\frac{d_{m}^{2}}{4\pi\epsilon_{0}}$ for electric dipoles and $g_{d}=\frac{d_{m}^{2}\mu_{0}}{4\pi}$ for magnetic dipoles, $d_{m}$ being the dipole moment, $\epsilon_{0}$ is the vacuum permittivity, and $\mu_{0}$ is the vacuum permeability. In general, the dipole-dipole interaction potential in 1D also includes a contact term owing to the transverse confinement, however that can be safely neglected for strong interaction strengths [102]. Unless otherwise stated, in the quench procedure we employ $g_{d}=100.0\to-100.0$. The parameter $\epsilon_{\alpha}$ is a short-range cutoff to avoid unphysical singularities arising at $x_{i}=x_{j}$. Note that $\epsilon_{\alpha}$ is varied according to the choice of $\alpha$. We tune $\alpha$ in the range $\left[0.5,4\right]$, which considers many different kinds of generalized long-range quantum systems. We choose the cut-off parameter such that the effective interaction $V_{\mathrm{eff}}$= $\int_{\mathcal{D}}{\frac{g_{d}}{x^{\alpha}+\epsilon_{\alpha}}\mathrm{d}x}=\int_{\mathcal{D}}{\delta(x)\mathrm{d}x}=1$, where the $\mathcal{D}=[-15,15]$ encompasses the entire simulation domain.
Before assessing the process of statistical relaxation for general $\alpha$, it is instructive to discuss the Fock space dynamics of the excited metastable sTG phase for a particular case. We select $\alpha$ = 3.0, and corresponding $\epsilon_{\alpha=3.0}$ $\approx 0.1015976$. To understand the dynamical fragmentation during TG to sTG quench, we perform a study with increasing orbital number $M$ and present results for the selected values of $M=5,10,15,20$.
IV.1 Violent fragmentation
Fig. 1 depicts the time evolution of the orbital occupation. The figure shows that initially $(t<0)$ the system is fragmented for all choices of orbital number. The occupation in the natural orbitals changes quite rapidly after the quenching procedure. The occupation in the orbitals with initial significant population decreases, while the one in the initially nearly empty orbitals increases. This trend continues until a saturation point, located approximately at time $t=0.5\bar{t}$ for all values of $M$, where the population of the orbitals coalesces around the value $1/M$. This represents an irregular behavior which we term violent fragmentation, and is the hallmark of a complete saturation of the probed Hilbert space of our quantum simulations. We remark that this type of violent fragmentation is exhibited — practically unchanged — for all values of $M$ up to $M=20$, which represents a computational upper threshold for our simulations given the current state-of-the-art. At the transition time the fully collapsed state always occupies the full available configuration space. The TG to sTG quench can therefore be described as a transition process from quantum to classical behavior of the long-range interacting bosons.
Figure 1: Orbital fragmentation dynamics for $N=5$ quenched bosons with $\alpha=3.0$, $g_{d}=100.0\to-100.0$ described with (a) $M=5$, (b) $M=10$, (c) $M=15$, (d) $M=20$ orbitals.
IV.2 Fock space delocalization
To further understand the quantum to classical behavior, we visualize the value of the coefficients $C_{\bar{n}}(t)$ after the quench in Fig. 2 We plot $|C_{\bar{n}}(t)|^{2}$ as a function of the index $n$ of the basis states for the final simulation time at $t=3.0\bar{t}$. The number of configurations in the available space for particular choice of $M$ is determined by $N_{\mathrm{conf}}$ by Eq.(4). Thus, $N_{\mathrm{conf}}$ increases exponentially from 126 for $M=5$ to over 40,000 for $M=20$. For each choice of $M$, the initial TG gas undergoes a violent delocalization process in which most of the configurations acquire a comparable weight. This behavior corresponds to the definition of classical gas, which occupies the whole available configuration space.
Figure 2: Configuration space population at time $t=3.0\bar{t}$. s measured from the amplitude squared of the many-body coefficients, for $N=5$ quenched bosons with $\alpha=3.0$, $g_{d}=100.0\to-100.0$ and increasing orbital number $M$. The orbital number is (a) $M=5$, (b) $M=10$, (c) $M=15$, (d) $M=20$.
IV.3 Unbounded Entropy measures
In the dynamical evolution of isolated quantum system, statistical relaxation and thermalization are established by some kind of equilibrium. For interacting many-body systems, statistical relaxation is also related to chaos in the energy spectra [103]. In particular, chaos in quantum interacting systems is defined by the statistics of the eigenstates. The onset of chaos is established by quadratic growth in entropy at very short time, then linear growth before saturation to a maximum entropy states. For time reversal and rotationally invariant systems, the maximum entropy is also determined by the predictions of a Gaussian orthogonal ensemble of random matrices(GOE) [104, 47, 105, 106].
Therefore, to further understand the violent dynamics and its impact on the relaxation process, in Fig 3 we plot the four different entropic quantities introduced in section III ($S_{C}(t)$, $S_{n}(t)$, $I_{C}(t)$ and $S_{C}^{N}(t)$) as measures of irregular dynamics. We present how the different measures approach the GOE prediction as determined by the number of orbitals used in the computation. For a GOE of random matrices, $I_{C}^{GOE}=D/3$ and $S_{C}^{GOE}=\ln 0.48D$, where $D\times D$ is the dimension of the random matrices [104]. To compare with GOE estimates, we set $D$ as equal to the number of configurations $N_{\mathrm{conf}}$ participating in the dynamics. For occupation entropy $S_{n}$, $D$ is set equal to the number of orbitals $M$ used in the computations. Thus $S_{n}^{GOE}$ = $-\sum_{i=1}^{M}\frac{1}{M}\ln(\frac{1}{M})=\ln(M)$. The GOE estimates are evaluated for different $M$ used in the MCTDH-B treatment.
Fig. 3 shows that statistical relaxation is achieved simultaneously with Fock space delocalization and violent dynamical fragmentation at time $t$=$0.5\bar{t}$. In particular, the results for increasing orbital number further support the thesis that the production of entropy is indeed unbounded. For example, for $M=15$, the number of configurations in the dynamics is $N_{conf}$ = 11’628. The corresponding GOE predictions are $S_{C}^{GOE}$ = 8.627, $S_{n}^{GOE}=2.708$, $I_{C}^{GOE}$= 3’876. We observe that for the prequench state $(t=0)$, all the measures are far enough from 00 indicating that the system initially is fragmented and delocalized. With time, all the entropy measures smoothly increase and then saturate. The saturation values for $M=15$ are $S_{C}^{\mathrm{sat}}\approx 8.94$, $S_{n}^{\mathrm{sat}}\approx 2.71$ and $I_{C}^{\mathrm{sat}}=5^{\prime}820$. These values are indeed close to GOE predictions, however a slight discrepancy exists because the interaction strength needed to exhibit TG limit is very large, albeit not infinite, which is the perquisite for GOE prediction. Repeating the same calculations with $M=20$ yields $N_{\mathrm{conf}}$=42’504, and the corresponding GOE predictions are $S_{C}^{GOE}$ = 9.923, $S_{n}^{GOE}=2.995$, $I_{C}^{GOE}$= 14’168. The corresponding saturation values in MCTDH-B computations are $S_{C}^{\mathrm{sat}}\approx 10.24$, $S_{n}^{\mathrm{sat}}\approx 2.99$ and $I_{C}^{\mathrm{sat}}\approx 21^{\prime}372$. The saturation close to GOE prediction indeed exhibits the hallmark of statistical relaxation. However, the impact of violent fragmentation is visible in the unbound growth of the entropy measures with the number of orbitals.
Figure 3: Dynamics for four different entropy measures for $N=5$ quenched bosons with $\alpha=3.0$, $g_{d}=100.0\to-100.0$ and increasing orbital number $M=5$ to $M=20$. (a) Coefficient entropy $S_{C}(t)$, (b) occupation entropy $S_{n}(t)$, (c) coefficient inverse participation ratio $I_{C}(t)$, (d) many-body coefficient entropy $S_{C}^{N}(t)$.
V Understanding of relaxation process
Understanding the relaxation process is an important unsolved problem in interacting quantum many-body systems. Theoretical concepts like the quantum ergodic theory or the eigenstate thermalization hypothesis (ETH) [106] infer possible relaxation processes described through generalized Gibbs ensembles. However, it is unclear on what time scale this occurs depending on the interaction range. It has been conjectured that several distinct time scales can arise in the dynamics of complex systems, although eventually they do settle to an equilibrium state in the long time dynamics [107].
An intriguing question to understand is the intermediate path from a relaxing state to a fully relaxed one. In this section, we thus analyze the many-body quantum dynamics over the entire range of the exponents $\alpha$ $\epsilon$ $\left[0.5,4\right]$. To understand the relaxation process we thoroughly discuss two typical cases: $\alpha=2.7$ (weakly long-range) and $\alpha=0.7$ (strongly long-range). Besides thoroughly investigating the dynamics of bosons, we also present calculations for spinless fermions to establish the universality of the fragmentation and relaxation processes. For both quantum statistics, we maintain the same protocol in the quench dynamics. We finally summarize the entire results in a parameter-time diagram which depicts the phase boundary between non-relaxed gas and relaxed gas.
V.1 Process of relaxation for bosons
We begin by probing the relaxation process for bosons. As before, we consider a sudden quench from $g_{d}=100.0\to-100.0$ with a fixed number of particles $N=5$, and concentrate on the results for $M=10$ orbitals.
First, we discuss the case for strong long-range interactions with $\alpha=0.7$ and measure all the four entropic quantities $S_{C}(t)$, $S_{n}(t)$, $I_{C}(t)$ and $S_{C}^{N}(t)$, which are plotted in Fig. 4. We observe a concurring estimate from all the four different measures that the relaxation happens around time $t\approx 0.6\bar{t}$. As expected, the saturation values approach the corresponding GOE predictions. At very small times, we observe a quadratic growth in $S_{c}(t)$ and $S_{C}^{N}(t)$ and both of them exhibit peculiar oscillations before reaching the saturation value. This behavior suggests a highly complex dynamics with several intermediate and rather unstable phases before the entropy eventually settles to the final state. $S_{n}$, instead, exhibits an almost linear increase at short times and smoothly settles to the GOE value. $S_{n}$ is more insensitive to irregularities in the dynamics as it is determined by the fragmentation of the many-body states. In fact, in the previous section we already saw that $S_{n}$ is the entropic quantity that best converges to the GOE result.
Figure 4: Dynamics for four different entropy measures for $N=5$ quenched, very long-range interacting bosons with $\alpha=0.7$, $g_{d}=100.0\to-100.0$, $M=10$. (a) Coefficient entropy $S_{C}(t)$, (b) occupation entropy $S_{n}(t)$, (c) coefficient inverse participation ratio $I_{C}(t)$, (d) many-body coefficient entropy $S_{C}^{N}(t)$.
The results for weak long-range interactions with $\alpha=2.7$ are presented in Fig. 5. The corresponding relaxation process is less complex than for strong long-range interactions. In the intermediate dynamics, the entropy measures exhibit a single but deeper minimum which is superseded by the eventual saturation to the GOE value. There is no clear signature of quasi equilibrium states for this case, and the time to reach the relaxed state is shorter ($t\approx 0.4\bar{t}$) than for strong long-range interactions.
Figure 5: Dynamics for four different entropy measures for $N=5$ quenched, nearly dipolar-interacting bosons with $\alpha=2.7$, $g_{d}=100.0\to-100.0$, $M=10$. (a) Coefficient entropy $S_{C}(t)$, (b) occupation entropy $S_{n}(t)$, (c) coefficient inverse participation ratio $I_{C}(t)$, (d) many-body coefficient entropy $S_{C}^{N}(t)$.
In Fig. 6, we summarize our results in the form of a parameter-time diagram across the interaction power $\alpha$ and the simulation time. Due to the closest agreement with GOE values for $S_{n}$ and its more regular dynamics that allows to determine the relaxation time more precisely, we will consider it as a good metric and use it as an order parameter in the parameter-time diagram. The diagram shows a clear separation between non-relaxed and relaxed states and we observe an exponential boundary between these two regimes. This suggests very long relaxation times in the limit $\alpha\to 0$, which cannot be reached with our current numerics due to the need of employing increasingly larger system sizes  [^3] We remark that the reduction of entropy exhibited by $\alpha=0.5$ at $t>0.6\bar{t}$ is likely due to self-interference of the long-range tails of the interacting particles. In our numerics, although we employ a harmonic confinement, the system is always defined with periodic boundary conditions. If the system size is not large enough, as it might have happened for the $\alpha=0.5$ results, self-interference patterns across the system boundaries will appear and modify the dynamics.. The parameter-time diagram has more distinct characteristics when $\alpha<1$, i.e. when the range is less than the dimension of the system. For larger values, instead, the exponential boundary gradually loses its structure and the relaxation time is almost independent of the choice of $\alpha$.
Figure 6: Occupation entropy dynamics as a function of interaction power $\alpha$ for $N=5$ quenched bosons with $g_{d}=100.0\to-100.0$ and $M=10$. The dashed line indicates the exponential boundary between thermalized and non-thermalized states.
V.2 Process of relaxation for fermions
In this section, we present similar many-body dynamics results for $N=5$ spinless fermions, maintaining the same prequench and postquench parameters as chosen for the bosonic system. The corresponding measures of entropy are plotted in Fig. 7 and Fig. 8 for $\alpha=0.7$ and $2.7$ respectively. In the strong long-ranged quench ($\alpha=0.7$), we observe highly modulated oscillatory structures in the dynamics of $S_{C}(t)$, $I_{C}(t)$ and $S_{C}^{N}(t)$. This makes it hard to determine the exact point of relaxation. However, we find that the occupation entropy exhibits a smoother behavior and saturates to the same GOE prediction observed for bosons, which establishes the universality of the many-body dynamics in the quantum quench to the strong interaction regime.
Figure 7: Dynamics for four different entropy measures for $N=5$ quenched, nearly dipolar-interacting fermions with $\alpha=0.7$, $g_{d}=100.0\to-100.0$, $M=10$. (a) Coefficient entropy $S_{C}(t)$, (b) occupation entropy $S_{n}(t)$, (c) coefficient inverse participation ratio $I_{C}(t)$, (d) many-body coefficient entropy $S_{C}^{N}(t)$.
For weak long-range interactions ($\alpha=2.7)$, oscillations with smaller amplitudes are observed, and the relaxation process becomes less complex. Eventual relaxation is observed in all the quantities and again $S_{n}$ agrees with the GOE prediction.
Figure 8: Dynamics for four different entropy measures for $N=5$ quenched, nearly dipolar-interacting fermions with $\alpha=2.7$, $g_{d}=100.0\to-100.0$, $M=10$. (a) Coefficient entropy $S_{C}(t)$, (b) occupation entropy $S_{n}(t)$, (c) coefficient inverse participation ratio $I_{C}(t)$, (d) many-body coefficient entropy $S_{C}^{N}(t)$.
The fermionic phase diagram as shown in Fig. 9 exhibits an exponential boundary between the relaxing and relaxed states already observed for bosons. However, the boundary is slightly shifted to larger times compared to the bosonic case when $\alpha<1$, whereas it remains rather unchanged for shorter range interactions.
Figure 9: Occupation entropy dynamics as a function of interaction power $\alpha$ for $N=5$ quenched fermions with $g_{d}=100.0\to-100.0$ and $M=10$. The dashed line indicates the exponential boundary between thermalized and non-thermalized states.
VI Control of relaxation
It is an established fact that a (near-)integrable system relaxes in at least two time scales related to the faster and slower relaxation processes [107]. Prethermalization occurs in the intermediate time scale before asymptotic relaxation to an equilibrium state. A necessary condition for thermalization is statistical relaxation in various observables to some kind of equlibrium. However, the observed dynamics in the TG to sTG quench is a highly nonintegrable phenomenon. The dynamics is so complex that different intermediate states may exist, but we can not detect them due to their extremely small lifetimes. It is thus impossible to establish the possibility of any prethermal state in the strongly quenched regime within the current computational limitations.
Nevertheless, another key question that can be addressed is whether it is possible to control the complex and extremely nontrivial pathways observed in the relaxation dynamics during the TG to sTG quench by probing regimes of weaker interactions. In the following, we thus present an extended analysis of the dynamics for quenches with $g_{d}=+100.0$ to $g_{d}=-80.0,-50.0,-30.0$ and $-10.0$. Again, we evaluate the measures of entropy dynamics for $N=5$ bosons for two distinct choices of $\alpha$. We present the results for $\alpha=0.7$ in Fig. 10 and for $\alpha=2.7$ in Fig. 11. For $\alpha=0.7$, we observe that the dynamics is strongly affected by decreasing the interactions to weaker values. The system exhibits more complex pathways and the relaxation process is severely hampered. In the same time scale, when $g_{d}$ is quenched from $+100.0$ to $-100.0$ and the system exhibits perfect relaxation after passing through many complex intermediate states, the $g_{d}$= $+100.0$ to $-10.0$ quench is unable to achieve a completely relaxed state.
It is known that a linear increase in entropy is the precursor of thermalization [109]. In Fig. 10 (b), we indeed observe that $S_{n}(t)$ exhibits a linear increase when the interaction strength is quenched to the same magnitude. Gradually weakening the quench, though, the linear increase starts to disappear and instead a plateau starts to develop. For $g_{d}=+100.0$ to $g_{d}=-10.0$ quench, the growth of $S_{n}(t)$ clearly exhibits three different time scales rather than a linear increase of the entropy. In the first interval (green background), there is an almost steady region. In the second interval (yellow background), $S_{n}(t)$ increases more rapidly and seems to follow logarithmic growth. In the third interval (red background), $S_{n}(t)$ becomes approximately constant and does not quite reach the GOE saturation value. These features indicate the possibility of generating a prethermal state by combining longer-ranged interactions with less violent quenches in generalized long-range systems.
Figure 10: Dynamics for four different entropy measures for $N=5$ quenched, nearly dipolar-interacting bosons with $\alpha=0.7$ and varying quench strength. (a) Coefficient entropy $S_{C}(t)$, (b) occupation entropy $S_{n}(t)$, (c) coefficient inverse participation ratio $I_{C}(t)$, (d) many-body coefficient entropy $S_{C}^{N}(t)$. For all curves $M=10$ orbitals were used. In panel (b), the three different background colors denote the three different regimes observed in the $g_{d}=-100.0\to+10$ quench.
The dynamics shown in Fig. 11 for shorter range interactions and smaller quench is distinctly different. The system remains mostly insensitive to the reduction in the quench strength. Some modulations in the amplitude are observed, but the basic conclusion remains unchanged: the system saturates to the GOE value in a single time scale, irrespective of the post-quench parameters. Fig 11(b) shows that $S_{n}(t)$ maintains its linearity at all times which guarantees thermal-like relaxation for all weaker quench processes. No intermediate plateau is formed. Notably, the saturation value is almost independent on the postquench parameter. We thus conclude that when $\alpha$ is smaller than the dimension of the system, the dynamics can be drastically modified by the strength of the quench protocol, whereas for shorter range interactions the dynamics remains insensitive to it.
Figure 11: Dynamics for four different entropy measures for $N=5$ quenched, nearly dipolar-interacting bosons with $\alpha=2.7$ and varying quench strength. (a) Coefficient entropy $S_{C}(t)$, (b) occupation entropy $S_{n}(t)$, (c) coefficient inverse participation ratio $I_{C}(t)$, (d) many-body coefficient entropy $S_{C}^{N}(t)$. For all curves $M=10$ orbitals were used.
VII Conclusions and Outlook
Understanding non-equilibrium dynamics and relaxation of quantum many-body systems is one of the major unsolved problems in quantum mechanics. The scarcity of appropriate theoretical tools to probe the relaxation dynamics makes the scenario even more challenging. When a generic isolated quantum system relaxes, the process of thermalization is not always known. There may be partial relaxation, or different stages of relaxation that can happen on different time scales leading to very complex nonequilibrium dynamics.
In this regard, thermalization is one of the most intriguing features and has been shown to emerge in various theoretical and experimental setups, although its physical origin is still not properly understood. Near-integrable quantum systems, which fail to thermalize on experimental time scales, can relax to some prethermal states described by the Gibbs ensemble. Existing numerical studies of the quench dynamics in the one-dimensional transverse field Ising model with long-range interactions $(\frac{1}{r^{\alpha}})$ exhibit signatures of prethermalization with subsequent dynamical phase transitions when $\alpha<d$ ($d$ being the dimension of the system).
The present work addresses the nonequilibrium quench dynamics in a long-range interacting system, albeit with a different quench protocol. We investigated a generalized long-range interacting Bose gas where the strong repulsive interactions are suddenly quenched to strong attractive values, following the seminal observation of the exotic sTG gas in Haller’s experiment [62]. We presented a numerical study of the many-body dynamics for a wide range of interactions for both bosonic and fermionic systems. The usage of ab initio many-body methods facilitates a complete understanding of the exotic sTG phase in terms of fragmentation, delocalization and entropy production. We observed that the dynamics exhibited by the TG to sTG quench reveals new insights into the structure of quantum state relaxation.
Although the system eventually relaxes to a maximal entropy state, it passes through very complex phases of violent fragmentation and chaotic delocalization. The entropy production is determined by the number of configurations in the Hilbert space, which are saturated for all the numbers of orbitals that can be employed with current computational capabilities. The observation of unbounded entropy and chaotic spreading of the sTG gas in the available Hilbert space strongly suggests a behavior akin to that of a classical gas. We presented different measures of entropy for our discussion. However, we chose the occupation entropy as an order parameter to produce a parameter-time diagram that illustrates the relaxation process as a function of interaction length, covering the entire range of $4.00.5<\alpha<$ 4.0. We found an exponential boundary separating the relaxing (quantum) from the relaxed (classical) phase. The same relaxation behavior is observed for spinless fermions, which establishes the universality of such chaotic dynamics in interacting quantum systems. We also studied a quench protocol to weaker attractions and observed signals of quasi-equilibrium for truly long-ranged interactions.
The theoretical findings obtained in this work for a broad range of decay exponents can be tested across various experimental systems such as trapped ions, dipolar systems, Rydberg states, and quantum gases in cavities. Each of these systems serves as a prototypical platform to examine quench dynamics in long-range physics. The typical scenarios involving confined bosons and fermions in the strongly interacting limit, as studied here, can also be implemented in various other spin models. These include the quantum Ising chain and the XYZ model, both of which allow complete control over the spin-spin interactions. Among these examples, trapped ions uniquely enable the realization of long-range interactions by adjusting the decay exponent. This adjustment is achieved by tuning the frequency of the detuned laser beat note and the trap frequency. Controlling the decay constant is fundamental to the ethos of quantum simulation. In the realm of quantum computation, trapped-ion systems are routinely examined, and qubit control is facilitated through the manipulation of long-range interactions. Therefore, the comprehensive analysis of the quench protocol presented in this manuscript can be effectively used as a quantum tool to manipulate dynamics and to explore quantum information tasks within the highly versatile platform of trapped-ion systems.
Acknowledgements
We gratefully acknowledge computation time on the ETH Zurich Euler cluster and at the High-Performance Computing Center Stuttgart (HLRS). We thank E. Bergholtz for useful discussions. This work is partly supported by the Swedish Research Council (2018-00313) and Knut and Alice Wallenberg Foundation (KAW) via the project Dynamic Quantum Matter (2019.0068).
Appendix A Initial states
In this section we show the density profile for the initial states used in the quench protocol. The initial state for $N=5$ bosons is presented in Fig. 12 (a). The initial state for $N=5$ fermions is presented in Fig. 12 (b). The real space density distribution is plotted as a function of the exponent $\alpha$. The interaction strength used is $g_{d}=100.0$ and the computation is performed with $M=10$ orbitals. We have verified that employing a higher number of orbitals does not change the density distribution.
Appendix B Stronger quench dynamics
In this section, we visualize and discuss the phase diagram of the relaxation process across different values of interaction exponent $\alpha$ for an even stronger quench than the one considered in the main text. This is shown in Fig. 13. Compared to the phase diagrams presented in the main text, we observe two different features. First, the relaxation process occurs at even shorter times than for the quench $g_{d}=100.0\to-100.0$, which is expected given the stronger interactions. Second, the boundary between the relaxing and relaxed states loses its exponential character almost completely, and instead appears to be better described by a linear function.
Figure 12: Real-space density distribution $\rho(x)$ as a function of the interaction power $alpha$ for the crystal states used as initial states in the quench procedures of the main text. (a) $N=5$ repulsive bosons, (b) $N=5$ repulsive fermions. For both panels an interaction strength $g_{d}=100.0$ and $M=10$ orbitals were used.
Figure 13: Occupation entropy dynamics as a function of interaction power $\alpha$ for $N=5$ quenched bosons with $g_{d}=100.0\to-200.0$ (i.e. a stronger quench than in the main text) and $M=10$.